OpenStudy (anonymous):

Use the substitution method to solve the system of linear equations. 2x-y=6 and 9x-4y=28. What is the solution?

OpenStudy (anonymous):

1) y = 2x - 6 2) 4y = 9x - 28 y = 9/4x - 7 1) = 2) 2x - 6 = 9/4x - 7 2x - 9/4x = -7 + 6 8/4x - 9/4x = -1 -1x = -1 x = 1 x = 1 --> 1) y = 2x - 6 y = 2(1) - 6 y = 2 - 6 y = -4

OpenStudy (anonymous):

So, the solution is (2,1)?

OpenStudy (anonymous):

(1,-4)

OpenStudy (anonymous):

Thank you! I have been struggling all semester!

OpenStudy (anonymous):

Your Welcome , Medal Please ?! (: